OpenStudy (anonymous):

find the quotient. (-15b^5+18b^3)/3b explain and check my answer.

OpenStudy (anonymous):

3b^8?

OpenStudy (anonymous):

@shinebrightlikeadimon

OpenStudy (anonymous):

do i solve it just as it looks?

OpenStudy (campbell_st):

nope... you need to split the numerator to and write the denominator with a power \[\frac{-15b^5}{3b^1} + \frac{18b^3}{3b^1}\]\\ so -15/3 = -5 18/3 = 6 now you need the index law for division, which is subtract the powers \[\frac{x^a}{x^b} = x^{a - b}\] so use this to simplify the b's hope it helps

OpenStudy (anonymous):

so then its im confused

OpenStudy (anonymous):

Okay you see how you told me you were trying to find the product?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

well your not it's quotient and it's different quotient is like if you were given the number like you 10 devided by 5 the quotient is 2

OpenStudy (anonymous):

yeah quotient thats what i was trying to come up with lol

OpenStudy (anonymous):

so then i dont solve it as it looks?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

like so do i just do what it shows?

OpenStudy (anonymous):

(-15b^5+18b^3)/3b

OpenStudy (anonymous):

so i do -15b^5 + 18b or do i mix it up?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

great now i got it wrong :d

OpenStudy (anonymous):

like what are the answer options

OpenStudy (anonymous):

there are none unfortunately

OpenStudy (anonymous):

oh im so sorry i was going to ask my dad and he's on the phone that's why i took so long

OpenStudy (anonymous):

but now its 49c^5-7c^2/-7c

OpenStudy (anonymous):

its okay lol :D no worries

OpenStudy (anonymous):

you need help with that one?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

im bad at math

OpenStudy (anonymous):

what type of math is this?

OpenStudy (anonymous):

algebra, but im retaking it online from freshaman year i have to make it up before next year cause ill be a senior lol

OpenStudy (anonymous):

yeah but is it algebra 1 or 2?

OpenStudy (anonymous):

1 lol

OpenStudy (anonymous):

my brain just cant learn big math equations good. its like i have to study it a while before i get it

OpenStudy (anonymous):

it's fine i did algebra 1 and i just recently finished geometry but idk i've never learned this so if you want you can just close this question post the question and tag someone who can help, i feel bad but i really don't remember doing this

OpenStudy (anonymous):

its fine :)

OpenStudy (anonymous):

Good luck!

OpenStudy (anonymous):

lol thanks :D

OpenStudy (anonymous):

np